### v1

# Classification of foliations via disintegration maps

Abstract.
In this paper, we present a novel approach for analyzing the relationship between the supports of conditional measures and their geometric arrangement in Wasserstein space via the disintegration map. Our method establishes criteria to determine when such conditional measures arise from a metric measure foliation. Additionally, we provide a concrete example demonstrating how this framework can be applied to study deformations (or perturbations) of disintegration-induced foliations.
1. Introduction
The disintegration of measures offers a powerful framework for analyzing local and global probabilistic structures by decomposing a measure into conditional probabilities. Roughly speaking, given $X,Y$ measurable spaces, $\mu$ a measure on $X$, and a measurable map $\pi:X\to Y$, a disintegration allows us to break $\mu$ into a family of conditional measures $\{\mu_{y}\}_{y\in Y}$ on $X$, depending measurably on $y$, such that $\mu$ can be reconstructed by integrating $\mu_{y}$ against $\pi_{*}\mu$, the push-forward of $\mu$ by $\pi$. This decomposition is crucial in several areas, such as Probability Theory, Ergodic Theory, and Geometric Measure Theory, where it provides a rigorous way to handle conditional distributions.
Disintegration theorems typically require topological assumptions to ensure existence and uniqueness of such decompositions, making them a powerful tool to define measures on fibered spaces or under projections, for instance. Although significant progress has been made in establishing the existence of conditional measures under several different hypothesis, most of the existing approaches neglect the geometric aspects of the underlying space.
In a recent paper [PR25], the second and third authors used the intrinsic structures of probability spaces to tackle the disintegration of measures from a perspective which takes into accounting statistical and geometrical properties of dynamical systems. In particular, it is introduced the notion of disintegration maps, which are essentially functions that assign to each point $y\in Y$ its corresponding conditional measure $\mu_{y}$, providing a structured way to analyze a disintegration. We emphasize that these objects can trace geodesics in the space of probability measures, which correspond to the solutions of a time-evolving transport problem, and exhibit certain rigidity phenomena: if one of the endpoints measure in the geodesic path is absolutely continuous, then so are all. Therefore, it becomes evident that certain geometric properties of the base space control the regularity of the conditional measures.
In this paper, we further explore disintegration maps by defining a notion of derivative for these objects. Our notion of derivative resembles a metric derivative, comparing in a appropriated way the distance between conditional measures and the distance between the supports of such measures. We remark that, in order to define this notion of derivative we do not require any differentiable structure on the base space. An interesting case arises in the context of Riemannian submersions, for example, since the distance between fibers equals the distance between base points. In this context, if each conditional measure has full support on its corresponding fiber, we show that the distance among conditional measures and the distance between base points are the same, and this isometry is captured by the derivative. We introduce a novel dense energy functional, similarly to what is done in geometric analysis. This notion is suited to quantify the relationship between these distances more accurately. We show that the minimization of this functional detect and characterize highly ordered geometric structures in the base space. More specifically, we show that the minimum of this energy functional corresponds to a metric measure foliation. That is, a foliation of the base space whose leaves are parallel to each other and where the distance between conditional measures on these leaves equals the distance between the leaves themselves. Moreover, this functional exhibits sensitivity to small deformations of the foliation, capturing fine geometric structures. We provide an explicit example in which we assess the sensitivity of the energy functional by inducing deformations in a metric measure foliation.
The main contributions of this paper are presented in Section 4, where we introduce a novel notion of derivative for disintegration maps (Definition 4.1). We then construct an energy functional designed to quantify the relationship between the Wasserstein distance of conditional measures and the distance between their supports (Definition 4.2). Our main result is:
Theorem 1.1.
The minimum of $\mathcal{E}(f)$, the energy of the disintegration map, is equals to one, if and only if, the supports of conditional measures form a metric measure foliation.
The precise statement shall be given in Theorem A.
The paper is organized as follows. Section 2 outlines the theoretical framework and necessary definitions. Section 3 reviews essential background on disintegration maps. Finally, Section 4 presents our core results, including a detailed example exploring the sensitivity of the energy functional to deformations of the leaves that support the conditional measures.
2. Theoretical framework
Let $X$, $Y$ be locally compact complete separable metric spaces. Consider $\pi:X\to Y$ a Borel map, and $\mu\in\mathcal{M}_{+}(X)$, where $\mathcal{M}_{+}(X)$ is the set of all finite positive Radon measures on $X$. Define $\nu:=\pi_{*}\mu$ in $\mathcal{M}_{+}(Y)$. Then, by [PR25, Theorem A] there exist measures $\mu_{y}\in\mathcal{M}_{+}(X)$, such that,
  1. (1)
    $y\mapsto\mu_{y}$ is a Borel map for $\nu$-almost every $y\in Y$;
  2. (2)
    $\mu_{y}$ is a probability on $X$ for $\nu$-almost every $y\in Y$;
  3. (3)
    $\mu_{y}$ is concentrated on $\pi^{-1}(y)$ for $\nu$-almost every $y\in Y$;
  4. (4)
    $\mu(A)=\int_{Y}\mu_{y}(A)~d\nu(y)$ for every Borel subset $A$ of $X$.
The family $\{\mu_{y}\}_{y\in Y}$ is called the disintegration of $\mu$ with respect to $\nu$.
in order to examine this family of probability measures, we endow $\mathscr{P}(X)$, the space of probability measures on $X$, with a metric structure. The framework of Wasserstein spaces provides a natural geometric setting for such investigations. The Wasserstein space of order $p$, for $p\in[1,\infty)$, is the set
$\mathscr{P}_{p}(X):=\Big{\{}\mu\in\mathscr{P}(X):\int d(x,\tilde{x})^{p}\mu(dx)<+\infty\Big{\}}$
$\tilde{x}\in X$ arbitrary, endowed with the Wasserstein distance
$W_{p}(\mu,\nu):=\Big{(}\inf\limits_{\gamma\in\Pi(\mu,\nu)}\int d(x_{1},x_{2})^{p}~d\gamma(x_{1},x_{2})\Big{)}^{\frac{1}{p}}.$
for $\mu$ and $\nu$ probability measures on $X$, where $\Pi(\mu,\nu)$ is the set of measures $\gamma\in\mathscr{P}(X\times X)$ with marginals $\mu$ and $\nu$.
Given a disintegration of $\mu$ with respect to $\nu$, we call disintegration map a function assigns to each $y\in Y$ a conditional measure $\mu_{y}\in\mathscr{P}_{p}(X)$.
Definition 2.1.
[PR25, Definition 4.1] Let $X$, $Y$ be locally compact complete separable metric spaces. Consider $\pi:X\to Y$ a Borel map, $\mu\in\mathcal{M}_{+}(X)$, and $\nu:=\pi_{*}\mu$. Given $\{\mu_{y}\}_{y\in Y}$ a disintegration of $\mu$ with respect to $\nu$, we define the disintegration map by
$f:D\subset Y$ $\to(\mathscr{P}_{p}(X),W_{p})$
$y$ $\mapsto\mu_{y},$
such that $\mu(A)=\int_{Y}f(y)(A)~d\nu(y)$, for every Borel subset $A$ of $X$. In order to make clear which measures are associated with the disintegration map, we say that “$f$ is a disintegration map of $\mu$ with respect to $\nu$”.
To give an example, consider $X=[0,1]\times[0,1]$ the unit square, $Y=[0,1]$, $\mu=\text{Leb}^{2}$ the two-dimensional Lebesgue measure on $X$, $\pi(x,y)=x$ the projection onto the first coordinate, and $\nu=\text{Leb}^{1}$ the one-dimensional Lebesgue measure on $Y$. The disintegration $\{\mu_{x}\}_{x\in[0,1]}$ of $\mu$ with respect to $\nu$ is given by:
$\mu_{x}=\text{Leb}^{1}\text{ on }\{x\}\times[0,1]\quad\text{(vertical line segment at }x\text{)}$
Explicitly, for any measurable $A\subseteq X$:
$\mu(A)=\int_{[0,1]}\mu_{x}(A)\,d\nu(x)=\int_{0}^{1}\text{Length}\big{(}A\cap(\{x\}\times[0,1])\big{)}\,dx$
where $\text{Length}(A\cap(\{x\}\times[0,1])):=\text{Leb}^{1}(y\in[0,1]~:~(x,y)\in A)$. In fact, note that the map $x\mapsto\mu_{x}(A)=\text{Length}(A\cap\{x\}\times[0,1])$ is measurable, for any $B\subseteq[0,1]$,
$\mu(A\cap\pi^{-1}(B))=\int_{B}\text{Length}\big{(}A\cap\{x\}\times[0,1]\big{)}\,dx$
and $\mu_{x}$ is supported on $\pi^{-1}(x)=\{x\}\times[0,1]$, satisfying the disintegration conditions.
Figure 1. Disintegration of $\text{Leb}^{2}$ with respect to $\text{Leb}^{1}$: the red line $\{x\}\times[0,1]$ carries $\mu_{x}=\text{Leb}^{1}$
In this case, the disintegration map is given by
$f:[0,1]$ $\to(\mathscr{P}_{p}([0,1]\times[0,1]),W_{p})$
$x$ $\mapsto\mu_{x},$
where $\mu_{x}$ is $\text{Leb}^{1}$ on $\{x\}\times[0,1]$. This is a basic example that fits into the broader framework of metric foliations, which is the primary focus of our study.
Let $(X,d)$ be a metric space, and $\mathcal{F}$ a partition of $X$ into closed subsets. We call $\mathcal{F}$ a foliation of $X$ and the elements of this partition are called leaves. If for every $F,F^{\prime}\in\mathcal{F}$ and every $x\in F$, $d(F,F^{\prime})=d(x,F^{\prime})$, where $d(F,F^{\prime})=\inf\{d(x,x^{\prime}):x\in F,x^{\prime}\in F^{\prime}\}$ and $d(x,F^{\prime})=d(\{x\},F^{\prime})$, then $\mathcal{F}$ is called a metric foliation. In case that each leaf is bounded, we say that $\mathcal{F}$ is bounded. Given a metric foliation $\mathcal{F}$ of $X$, define the equivalence relation:
$x\sim x^{\prime}~\iff~\exists~F\in\mathcal{F}~\text{such that}~x,x^{\prime}\in F.$ (1)
Consider the $X^{*}:=X/\sim$ the set of equivalence classes under (1) and the projection $\pi:X\to X^{*}$ onto $X^{*}$. We call $X^{*}$ the quotient space and $\pi$ the quotient map. Define a distance function $d^{*}$ on $X^{*}$ as
$d^{*}(y,y^{\prime}):=d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))$ (2)
for $y,y^{\prime}\in X^{*}$. Note that $\pi$ is a submetry: $\pi(B(x,r))=B(\pi(x),r)$, where $B(x,r)$ is a ball centred at $x$ with radius $r$. In fact,
$B(\pi(x),r)=$ $~\{y\in X^{*}:d^{*}(y,\pi(x))<r\}$
$=$ $~\{y\in X^{*}:d(\pi^{-1}(y),\pi^{-1}(\pi(x)))<r\}$
$=$ $~\pi(B(x,r)).$
Therefore, $\pi$ is 1-Lipschitz. In this notation, we define a metric measure foliation.
Definition 2.2.
Let $X$ be locally compact complete separable metric space, and $\mathcal{F}$ be a metric foliation of $X$. Consider $\mu\in\mathscr{M}_{+}(X)$, the quotient space $X^{*}$, and the quotient map $\pi:X\to X^{*}$. We call $\mathcal{F}$ a $p$-metric measure foliation, $p\in[1,\infty)$, if $\pi_{*}\mu$ is locally finite Borel measure on $X^{*}$, and
$W_{p}(\mu_{y},\mu_{y^{\prime}})=d^{*}(y,y^{\prime})$ (3)
for any $y,y^{\prime}\in Y$, where $\{\mu_{y}\}_{y\in Y}$ is a disintegration of $\mu$ with respect to $\pi_{*}\mu$.
Remark 2.3.
The general definition for a metric measure foliation requires $p=2$ and (3) only almost everywhere (see [GKMS18]). However, our definition is more restrictive, requiring (3) everywhere.
Note that, in this case, the disintegration map is an isometry. This is the most important characteristic of metric measure foliation for us. Although the definition holds for $p\in[1,\infty)$, our case of interest is $p=2$, since $\mathscr{P}_{2}(X)$ has a rich geometric structure [GRS25]. Moreover, if $\mathcal{F}$ is a 2-metric measure foliation, then
$W_{q}(\mu_{y},\mu_{y^{\prime}})=d^{*}(y,y^{\prime})=d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))$
for any $q\in[1,\infty)$ and any $y,y^{\prime}\in X^{*}$, where $\{\mu_{y}\}_{y\in Y}$ is the disintegration of $\mu$ with respect to $\pi_{*}\mu$ [Kaz22, Proposition 3.11]. Henceforth, the term “metric measure foliation” will refer to what we have defined as a “2-metric measure foliation”.
A very important example of metric measure foliation is related to the action of isometry group [GR23]. Let $(X,d,\mu)$ be a metric measure space and $G$ a compact topological group. Let
$G\times X\ni(g,x)\mapsto gx\in X$
be an isometric action of $G$ on $X$. Suppose this action is metric measure isomorphic. That is, for every $g\in G$ the map $X\ni x\mapsto gx\in X$ is an isometry preserving the measure $\mu$. Consider $[x]$ the $G$-orbit of a point $x\in X$ and the quotient space $X/G$ endowed with the distance
$d_{X/G}([x],[x^{\prime}])=\inf_{g,g^{\prime}\in G}d(gx,g^{\prime}x^{\prime}).$
Consider $\pi:X\to X/G$ the projection map, that is, $\pi$ is given by $x\mapsto[x]$. The family $\mathcal{F}:=\{\pi^{-1}(y):y\in X/G\}$ is a metric measure foliation on $X$.
Other interesting examples arise from Riemannian submersions of weighted Riemannian manifolds [GKMS18], for instance.
3. Preliminaries on disintegration maps
In this section, we discuss key properties of disintegration maps, as introduced in Definition 2.1 and developed in [PR25]. To start our discussion, suppose $X=Y=\mathbb{R}$, and let $\gamma$ be a probability measure supported on the graph of a measurable function $\pi:X\to Y$, i.e., $\gamma=(\mathrm{id},\pi)_{*}\mu$, where $\mu$ its a probability measure on $X$ and $(\mathrm{id},\pi)_{*}\mu$ is the pushforward of $\mu$ under the map $x\mapsto(x,\pi(x))$. Then, the disintegration of $\gamma$ with respect to $\mu$ is given by the family of Dirac measures:
$\gamma_{x}=\delta_{\pi(x)}.$
In fact, for any measurable $A\subseteq X\times Y$, we have
$\gamma(A)=\mu(\{x\in X:(x,\pi(x))\in A\}).$
By definition of the disintegration
$\gamma(A)=\int_{X}\gamma_{x}(A_{x})\,d\mu(x)=\int_{X}\delta_{\pi(x)}(A_{x})\,d\mu(x).$
where $A_{x}=\{y\in Y:(x,y)\in A\}$. Then,
$\delta_{\pi(x)}(A_{x})=\mathbbm{1}_{A}(x,\pi(x)),$
and thus:
$\gamma(A)=\int_{X}\mathbbm{1}_{A}(x,\pi(x))\,d\mu(x),$
which matches the definition of $\gamma$. This shows that when $\gamma$ is supported on the graph of a function $\pi$, its disintegration is given by $\gamma_{x}=\delta_{\pi(x)}$. This is a fundamental example in Optimal Transport and conditional probability, where deterministic couplings lead to Dirac disintegrations. Observe that when $\pi=\mathrm{id}$, we have a natural isometric immersion of $X$ into $\mathscr{P}_{p}(X)$, namely the disintegration map $x\mapsto\delta_{x}$.
In general, we aim to investigate under what conditions geometric information about $X$ can be derived via disintegration — specifically, through the properties of disintegration maps. A key starting point for this analysis is understanding the continuity of these maps. When referring to continuity of the disintegration map, we mean continuity with respect to weak convergence on $\mathscr{P}_{p}(X)$. In [PR25] some results are presented regarding the continuity of disintegration maps, namely:
  1. (1)
    [PR25, Propostion 5.2] Let $X$, $Y$ be locally compact complete separable metric spaces, and consider $\mu\in\mathcal{M}_{+}(X)$. If $\pi:X\to Y$ is a Borel map such that $\nu:=\pi_{*}\mu$ is a Borel measure, then, the disintegration map $f:Y\to(\mathscr{P}_{p}(X),W_{p})$ of $\mu$ with respect to $\nu$ is nearly continuous, for $p\in[1,\infty)$. A map $f$ on an metric space $(Y,\nu)$ is called nearly continuous if, for each $\varepsilon>0$ there exist $\mathcal{K}\subset Y$ closed with $\nu(Y\backslash\mathcal{K})<\varepsilon$ such that $f$ restricted to $\mathcal{K}$ is continuous.
  2. (2)
    [PR25, Propostion 5.4] Let $X$, $Y$ be locally compact complete separable metric spaces, and consider $\mu\in\mathcal{M}_{+}(X)$. If the Borel map $\pi:X\to Y$ is bijective and $\pi^{-1}$ continuous, then the disintegration map $f$ of $\mu$ with respect to $\pi_{*}\mu$ is continuous.
  3. (3)
    [PR25, Propostion 5.8] Let $X$ be locally compact complete separable metric space. Considering a metric foliation $\mathcal{F}$ of $X$, the quotient space $X^{*}$, and the quotient map $\pi:X\to X^{*}$, if there exists a metric measure foliation of $X$, then the disintegration map of $\mu$ with respect to $\nu:=\pi_{*}\mu$ is continuous.
The continuity of disintegration maps enables us to relate paths in the chosen indexing metric space to corresponding paths of measures in Wasserstein space. For instance, in the case of the disintegration map $x\mapsto\delta_{x}$, there is a direct correspondence between a path in $X$ and the induced path of Dirac measures. In the context of a metric measure foliation, this relationship becomes more nuanced. Here, transversal paths (those crossing the leaves) must satisfy an additional condition concerning the distance between the supports of the conditional measures along the path — a consequence of the parallelism of the leaves.
In the next section we present an approach that allows us to study the relation between the relative position of the elements of the family of supports and the relative position of the conditional measures in the Wasserstein space, via a disintegration map.
4. On foliations and disintegration maps
Motivated by the concept of the metric derivative and the interplay between the relative positions of supports and conditional measures, we introduce a notion of derivative for disintegration maps to quantify how these measures evolve along paths.
Definition 4.1.
Let $X,Y$ be locally compact complete separable metric spaces. Consider a measure $\mu\in\mathcal{M}_{+}(X)$, $\pi:X\to Y$ a Borel map, $\nu=\pi_{*}\mu$, and $f$ the disintegration map of $\mu$ with respect to $\nu$. We define a notion of derivative of $f$ by
$|\nabla f(y)|_{p}:=\lim\limits_{\varepsilon\to 0}\sup_{y^{\prime},y^{\prime\prime}\in B_{\varepsilon}(y)}\frac{W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})}{d(\pi^{-1}(y^{\prime}),\pi^{-1}(y^{\prime\prime}))}$
for $p\in[1,\infty)$, where $\mu_{y^{\prime}}=f(y^{\prime})$ and $\mu_{y^{\prime\prime}}=f(y^{\prime\prime})$ are conditional measures and $B_{\varepsilon}(y)$ is a ball centered in $y\in Y$ with radius $\varepsilon$.
Let $d_{Y}(y^{\prime},y^{\prime\prime})=d(\pi^{-1}(y^{\prime}),\pi^{-1}(y^{\prime\prime}))$ denote the distance between the fibers over $y^{\prime}$ and $y^{\prime\prime}$. In the case of a metric measure foliation, the $p$-Wasserstein distance between conditional measures satisfies $W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})=d_{Y}(y^{\prime},y^{\prime\prime})$, which implies $|\nabla f(y)|_{p}=1$. This derivative concept thus serves as a tool to characterize when $\{\pi^{-1}(y)\}$ forms a metric measure foliation. To further analyze the connection between the distances of conditional measure supports and their Wasserstein distance, paralleling techniques in geometric analysis, we introduce a dense energy functional designed to quantify this relationship.
Definition 4.2.
Let $X,Y$ be locally compact complete separable metric spaces. Consider a measure $\mu\in\mathcal{M}_{+}(X)$, a Borel map $\pi:X\to Y$, and the disintegration map $f$ of $\mu$ with respect to $\nu=\pi_{*}\mu$. The $p$-energy of $f$ is given by
$\mathcal{E}_{p}(f):=\|\nabla f\|_{\infty,p}=\sup_{y\in Y}|\nabla f(y)|_{p}.$
Note that $|\nabla f|_{p}\geq 1$ everywhere. In fact, since $\mu_{y^{\prime}}$ and $\mu_{y^{\prime\prime}}$ are supported on $\pi^{-1}(y^{\prime})$ and $\pi^{-1}(y^{\prime\prime})$, respectively, we have
$W_{p}^{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}}):=$ $\inf\limits_{\gamma\in\Pi(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})}\int d(x_{1},x_{2})^{p}~d\gamma(x_{1},x_{2})$
$\geq$ $~d(\pi^{-1}(y^{\prime}),\pi^{-1}(y^{\prime\prime}))^{p}.$
Moreover, note that $\min_{f}\mathcal{E}_{p}(f)=1$ if, and only if, $\|\nabla f\|_{\infty,p}=1$, which in turn hold true if, and only if $|\nabla f|_{p}=1$. In this case we have an isometry $W_{p}=d_{Y}$. Under some further hypotheses, we can relate the functional $\mathcal{E}_{p}(f)$ to the metric measure foliation case.
Theorem A.
Let $(X,d)$ be a locally compact complete separable geodesic space, $Y$ a compact complete separable metric space, and $\pi:X\to Y$ a Borel map. Given $\mu\in\mathcal{M}_{+}(X)$, let $f_{\mu,\pi}$ be a disintegration of $\mu$ with respect to $\nu:=\pi_{*}\mu$. Suppose that the conditional measures $\mu_{y}$ have full support, i.e., $\operatorname{supp}(\mu_{y})=\pi^{-1}(y)$. Then, $\mathcal{E}_{p}(f_{\mu,\pi})=1$ if, and only if, $\{\pi^{-1}(y)\}$ defines a $p$-metric measure foliation on $X$. Therefore, the energy-minimizing disintegration maps correspond to a metric measure foliation.
Proof.
[…]
$u(y^{\prime},y^{\prime\prime}):=\begin{cases}\frac{W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})}{d_{Y}(y^{\prime},y^{\prime\prime})}&~\text{,if}~~y^{\prime}\neq y^{\prime\prime}\\ 1&~\text{,if}~~y^{\prime}=y^{\prime\prime}\end{cases}$
Note that $u\geq 1$. By the hypothesis,
$\lim\limits_{\varepsilon\to 0}\sup\limits_{y^{\prime},y^{\prime\prime}\in B_{\varepsilon}(y)}u(y^{\prime},y^{\prime\prime})=u(y,y)$
for every $y\in Y$. Then, $u$ is continuous at $(y,y)$ for every $y\in Y$. Moreover, $u$ is continuous everywhere else. In fact, since $\min_{f_{\mu,\pi}}\mathcal{E}_{p}(f_{\mu,\pi})=1$, we have $|\nabla f(y)|p=1$ for every $y\in Y$, and the disintegration map $f$ is continuous. Thus, it implies the continuity of $u$ everywhere else.
Since $Y$ compact, $Y\times Y$ is compact and by compactness, we have uniform continuity, that is, for all $\varepsilon>0$ there exists $\delta>0$ such that,
$d_{Y\times Y}((y,y^{\prime}),(\tilde{y},\tilde{y}^{\prime}))\leq\delta$
implies $|u(y,y^{\prime})-u(\tilde{y},\tilde{y}^{\prime})|\leq\varepsilon$, for every ($y~)(y,y^{\prime}$),($\tilde$$\tilde{y}^{\prime}$), where
[…]
Note that $|u(y,y)-u(\tilde{y},\tilde{y}^{\prime})|=u(\tilde{y},\tilde{y}^{\prime})-1$. Then, setting $y=\tilde{y}$, and if $d_{Y}((\tilde{y},\tilde{y}^{\prime}),(\tilde{y},\tilde{y}^{\prime}))\leq\delta$, we get $d_{Y\times Y}((y,y),(\tilde{y},\tilde{y}^{\prime}))\leq\delta$, Thus,
$|u(y,y)-u(\tilde{y},\tilde{y}^{\prime})|=u(\tilde{y},\tilde{y}^{\prime})-1\leq\varepsilon,$
[…]
$\sum_{k=0}^{n-1}d(x_{k},x_{k+1})=d(x,x^{\prime}).$
Then,
[…]
$\leq$ $\sum_{k=0}^{n-1}d(\pi^{-1}(\pi(x_{k})),\pi^{-1}(\pi(x_{k+1})))$
$\leq$ $(1+\varepsilon)\sum_{k=0}^{n-1}d(\pi^{-1}(\pi(x_{k})),\pi^{-1}(\pi(x_{k+1})))$
$\leq$ $(1+\varepsilon)\sum_{k=0}^{n-1}d(x_{k},x_{k+1})$
$=$ $(1+\varepsilon)d(x,x^{\prime})$
[…]
Let $\gamma$ be the optimal transport plan from $\mu_{y}$ to $\mu_{y^{\prime}}$, and consider, for all $(x,x^{\prime})\in\operatorname{supp}(\gamma)$, $\varphi(x)=d(x,\pi^{-1}(y^{\prime}))>d(x,x^{\prime})$. Then,
$\int\varphi(x)d\gamma=\int\varphi(x)d\mu_{y}=\int d(x,\pi^{-1}(y^{\prime}))d\mu_{y}>\int d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))d\mu_{y}$
that is a contradiction, since
$W_{1}(\mu_{\tilde{y}},\mu_{\tilde{y}^{\prime}}):=\inf\limits_{\gamma\in\Pi(\mu_{\tilde{y}},\mu_{\tilde{y}^{\prime}})}\int d(x_{1},x_{2})~d\gamma(x_{1},x_{2}).$
An analogous argument applies to the case $p\in(1,\infty)$.
Remark 4.3.
The conclusion of Theorem A does not hold if full support is not assumed. In fact, consider $X=[-1,1]\times\mathbb{R}$, $Y=[0,1]$, $\pi:X\to Y$ such that $\pi^{-1}(y)$ is a ellipse given by $x_{1}^{2}+\lambda^{2}x_{2}^{2}=y^{2}$, with major radius $y$. Consider a measure $\mu\in\mathcal{M}_{+}(X)$ such that the disintegration of $\mu$ with respect to $\nu$ is $\mu_{y}=\delta_{(0,\frac{y}{\lambda})}$. In this case,
$d_{Y}(\pi^{-1}(y),\pi^{-1}(y^{\prime}))=\left|\frac{y}{\lambda}-\frac{y^{\prime}}{\lambda}\right|=W_{p}(\mu_{y},\mu_{y^{\prime}})$
for every $y,y^{\prime}\in Y$. Therefore, we have isometry, but $\{\pi^{-1}(y)\}$ does not define a metric foliation. A graphical representation of the case $\lambda=1.5.$ is provided in Figure 2.
Figure 2. Illustration for the case in which the measures $\mu_{y}$ do not have full support: there is isometry between $W_{p}$ and $d_{Y}$, but $\{\pi^{-1}(y)\}$ (the ellipses) does not define a metric foliation.
Note that in the previous example, if $\lambda=1$, that is, the preimages are circles in $X$, then $\{\pi^{-1}(y)\}$ forms a metric foliation. In the context of Theorem A, we aim to demonstrate that the $p$-energy exhibits sensitivity to small perturbations of the foliation. To analyze this, we first evaluate the $1$-energy $\mathcal{E}_{1}$ for a metric measure foliation on $\mathbbm{R}^{2}$ with concentric circular leaves. We then generalize this analysis to cases where the leaves are slightly deformed into ellipses.
Example 4.4.
Consider $X=\{(x_{1},x_{2})\in\mathbb{R}^{2}~:~x_{1}^{2}+x_{2}^{2}\leq R^{2}\}$, $R>0$, $\mu\in\mathscr{P}(X)$ with full support and uniformly distributed, and $Y=(0,R]$. Consider a foliation of $X$ into circles, that is, consider a foliation $\mathcal{F}$ of $X$ such that $F_{y}=\{(x_{1},x_{2})\in\mathbb{R}^{2}~:~x_{1}^{2}+x_{2}^{2}=y^{2}\}$, for each $y\in Y$, and $\pi:X\to Y$ such that $\pi((x_{1},x_{2}))=\sqrt{x_{1}^{2}+x_{2}^{2}}=y$. To simplify the analysis, consider this setup in polar coordinates: consider $x_{1}=r\cos(\theta)$, $x_{2}=r\sin(\theta)$. Using Cartesian coordinates, an infinitesimal area element can be calculated as
$dA=dx_{1}~dx_{2}.$
To convert it to polar coordinates, the Jacobian determinant $J$ of the coordinate conversion formula has to be considered. In this case:
$J(r,\theta)=\begin{vmatrix}\frac{\partial x_{1}}{\partial r}&\frac{\partial x_{1}}{\partial\theta}\\ \frac{\partial x_{2}}{\partial r}&\frac{\partial x_{2}}{\partial\theta}\end{vmatrix}=\begin{vmatrix}\cos(\theta)&-r\sin(\theta)\\ \sin(\theta)&r\cos(\theta)\end{vmatrix}=r.$
Then,
$dA=r~dr~d\theta.$
Since $\mu$ is a uniformly distributed probability measure on $X$, we can write:
$d\mu=\frac{dA}{A}=\frac{r~dr~d\theta}{\pi R^{2}}.$
Define $\nu:=\pi_{*}\mu$. Note that, give a open interval $(a,b)\subset(0,R]$, $\nu((a,b))$ is the area between circles of radii $a$ and $b$, normalized by $A$:
$\nu((a,b))=\frac{\pi b^{2}-\pi a^{2}}{\pi R^{2}}=\frac{b^{2}-a^{2}}{R^{2}}.$
In this way,
$d\nu(r)=\frac{2r}{R^{2}}~dr.$
Let $\{\mu_{y}\}_{y\in Y}$ be the disintegration of $\mu$ with respect to $\nu$, such that each $\mu_{y}$ is a uniformly distributed probability measure on $F_{y}$. Due to the symmetry of the circle, the arc length is proportional to $\theta$, and
$d\mu_{y}(r,\theta)=\frac{1}{2\pi}~\delta_{y}~d\theta.$
Note that
$\int_{Y}\mu_{y}~\nu=\int_{0}^{R}\int_{0}^{2\pi}\Big{(}\frac{1}{2\pi}\delta_{y}d\theta\Big{)}\frac{2r}{R^{2}}~dr=\frac{1}{\pi R^{2}}\iint r~dr~d\theta$
so we recover the measure $\mu$ by integrating the family $\mu_{y}$ with respect to $\nu$. Moreover, $d_{Y}(y,y^{\prime}):=d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))=|y-y^{\prime}|$. For simplicity, consider $W_{1}$. One way to write the 1-Wasserstein distance is
$W_{1}(\mu_{y},\mu_{y^{\prime}})=\inf_{T}\int d(x,T(x))~d\mu_{y}$
and since we have measures uniformly distributed and symmetry in the leaves, the function $T$ that minimizes this problem is such that, for every $x\in F_{y}$, $T(x)$ is a point in $F_{y^{\prime}}$ such that $d(x,T(x))=|y-y^{\prime}|$, that is, the transport occurs radially, orthogonally to the leaves. Therefore,
$W_{1}(\mu_{y},\mu_{y^{\prime}})=\int_{F_{y}}|y-y^{\prime}|~d\mu_{y}=|y-y^{\prime}|.$
Then, the disintegration map $\tilde{f}$ of $\mu$ with respect to $\nu$ is a isometry, and $\mathcal{E}_{1}(\tilde{f})=1$.
Similarly, consider the case $X=\{(x_{1},x_{2})\in\mathbb{R}^{2}~:~x_{1}^{2}+\lambda^{2}x_{2}^{2}\leq R^{2}\}$, with $R>0$ and $\lambda=1+\varepsilon$, $\varepsilon$ positive and sufficiently small, $\mu\in\mathscr{P}(X)$ with full support and uniformly distributed, $Y=(0,R]$, and the foliation of $X$ is given into ellipses, that is, consider a foliation $\mathcal{F}$ of $X$ such that $F_{y}=\{(x_{1},x_{2})\in\mathbb{R}^{2}~:~x_{1}^{2}+\lambda^{2}x_{2}^{2}=y^{2}\}$, for each $y\in Y$, and $\pi:X\to Y$ such that $\pi((x_{1},x_{2}))=\sqrt{x_{1}^{2}+\lambda^{2}x_{2}^{2}}=y$. Converting into polar coordinates: consider $x_{1}=r\cos(\theta)$, $x_{2}=\frac{r}{\lambda}\sin(\theta)$, and
$J(r,\theta)=\begin{vmatrix}\frac{\partial x_{1}}{\partial r}&\frac{\partial x_{1}}{\partial\theta}\\ \frac{\partial x_{2}}{\partial r}&\frac{\partial x_{2}}{\partial\theta}\end{vmatrix}=\begin{vmatrix}\cos(\theta)&-r\sin(\theta)\\ \frac{1}{\lambda}\sin(\theta)&\frac{r}{\lambda}\cos(\theta)\end{vmatrix}=\frac{r}{\lambda}.$
Then,
$dA=\frac{r}{\lambda}~dr~d\theta.$
Since $\mu$ is a uniformly distributed probability measure on $X$, we can write:
$d\mu=\frac{dA}{A}=\frac{\frac{r}{\lambda}~dr~d\theta}{\frac{\pi R^{2}}{\lambda}}=\frac{r~dr~d\theta}{\pi R^{2}}.$
Define $\nu:=\pi_{*}\mu$. Note that, give a open interval $(a,b)\subset(0,R]$, $\nu((a,b))$ is the area between ellipses of major axis $a$ and $b$, normalized by $A$:
$\nu((a,b))=\frac{\frac{\pi b^{2}}{\lambda}-\frac{\pi a^{2}}{\lambda}}{\frac{\pi R^{2}}{\lambda}}=\frac{b^{2}-a^{2}}{R^{2}}.$
In this way,
$d\nu(r)=\frac{2r}{R^{2}}~dr.$
Let $\{\mu_{y}\}_{y\in Y}$ be the disintegration of $\mu$ with respect to $\nu$, such that each $\mu_{y}$ is a uniformly distributed probability measure on $F_{y}$. However, note that in this case the relation arc length and $\theta$ is no longer linear (see Figure 3).
Figure 3. Normalized arc length as a function of $\theta$ for $\lambda$ equals to $1,1.1,1.5$ and $2$.
Therefore, the measures $\mu_{y}$ have a different dependence on $\theta$:
$d\mu_{y}(r,\theta)=\frac{\frac{dL_{y}(\theta)}{d\theta}}{\mathcal{L}_{y}}~\delta_{y}~d\theta$
where
$L_{y}(\theta)=\int_{0}^{\theta}y~\sqrt{1-(1-\lambda^{-2})\cos^{2}(\tilde{\theta})}~d\tilde{\theta}$
is the arc length of the ellipse $x_{1}^{2}+\lambda^{2}x_{2}^{2}=y^{2}$ for angle between 0 and $\theta$, and $\mathcal{L}_{y}=L_{y}(2\pi)$. By the Fundamental Theorem of Calculus
$d\mu_{y}(r,\theta)=\frac{y~\sqrt{1-(1-\lambda^{-2})\cos^{2}(\theta)}}{\mathcal{L}_{y}}~\delta_{y}~d\theta.$
In this case, $d_{Y}(y,y^{\prime})=d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))=\frac{1}{\lambda}|y-y^{\prime}|$ and, to calculate the $1$-Wasserstein distance between $\mu_{y}$ and $\mu_{y^{\prime}}$, we still have the radial transport, although $d(x,T(x))$ depends of $\theta$: for $x=y\cos(\theta)$, $d(x,T(x))=R_{y^{\prime}}(\theta)-R_{y}(\theta)$, where
$R_{y}(\theta)=\frac{y}{\sqrt{\lambda^{2}-(\lambda^{2}-1)\cos^{2}(\theta)}}$
and
$R_{y^{\prime}}(\theta)=\frac{y^{\prime}}{\sqrt{\lambda^{2}-(\lambda^{2}-1)\cos^{2}(\theta)}}.$
Then,
$W_{1}(\mu_{y},\mu_{y^{\prime}})=$ $\int R_{y^{\prime}}(\theta)-R_{y}(\theta)~d\mu_{y}$
$=$ $\int\frac{y^{\prime}-y}{\sqrt{\lambda^{2}-(\lambda^{2}-1)\cos^{2}(\theta)}}\frac{y~\sqrt{1-(1-\lambda^{-2})\cos^{2}(\theta)}}{\mathcal{L}_{y}}~\delta_{y}~d\theta$
$=$ $\frac{(y^{\prime}-y)~y}{\lambda~\mathcal{L}_{y}}~\int_{0}^{2\pi}~d\theta$
$=$ $\frac{2\pi~(y^{\prime}-y)~y}{\lambda~\mathcal{L}_{y}}.$
Therefore, considering $\hat{f}_{\lambda}$ the disintegration map of $\mu$ with respect to $\nu$, we have
$\mathcal{E}_{1}(\hat{f}_{\lambda})=\frac{2\pi y}{\mathcal{L}_{y}},$
that is, the energy is the ratio between the length of the circle with radius $y$ and the length of the ellipse with major axis equal to $y$. Table 1 displays the values of $\mathcal{L}_{y}$ for $\lambda=1$ (circle), $\lambda=1,001$, $\lambda=1,01$, $\lambda=1,1$, $\lambda=1,5$ and $\lambda=2$. Table 2 summarizes the computed values of $\mathcal{E}_{1}(\hat{f}_{\lambda})$ corresponding to the cases in Table 1. Observe that $\mathcal{E}_{1}(\hat{f}_{\lambda})$ is independent of $y$.
Table 1. $\mathcal{L}_{y}$ for different values of $\lambda$
y $\lambda=1$ $\lambda=1,001$ $\lambda=1,01$ $\lambda=1,1$ $\lambda=1,5$ $\lambda=2$
11 $6,28318531$ $6,28004725$ $6,25211912$ $6,00098645$ $5,28847986$ $4,84422411$
Table 2. Values of $\mathcal{E}_{1}(\hat{f}_{\lambda})$
$\lambda=1$ $\lambda=1,001$ $\lambda=1,01$ $\lambda=1,1$ $\lambda=1,5$ $\lambda=2$
11 $1,000499687$ $1,004968906$ $1,047025412$ $1,188089106$ $1,297046785$
Figure 4 depicts the dependence of $\mathcal{E}_{1}(\hat{f}_{\lambda})$ on $\lambda$. As shown, $\mathcal{E}_{1}$ grows smoothly with increasing foliation perturbation.
Figure 4. $\mathcal{E}_{1}(\hat{f}_{\lambda})$ as a function of $\lambda$
This example highlights the versatility of our approach: beyond classifying foliations on $X$ via disintegration maps, it also provides a framework for analyzing structural perturbations in foliations. Notably, the concepts developed here have potential applications across dynamical systems, machine learning, and stochastic analysis.
Acknowledgements
This work was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001. C. S. R. has been partially supported by São Paulo Research Foundation (FAPESP): grant #2016/00332-1, grant #2018/13481-0, and grant #2020/04426-6. The opinions, hypotheses and conclusions or recommendations expressed in this work are the responsibility of the authors and do not necessarily reflect the views of FAPESP.
C. S. R. would like to acknowledge support from the Max Planck Society, Germany, through the award of a Max Planck Partner Group for Geometry and Probability in Dynamical Systems. R. P. would like to acknowledge support from the Max Planck Institute for Mathematics in the Sciences and from CAPES.